Three blocks of masses 1kg,4kg,2kg are placed on a frictionless horizontal plane surface.two forces 120N and 50N are acting on the surface.120 N acts on the 1kg mass particle and 50 N acts in the opposite direction on 2kg masses.Find accelaration of the system.

Grade:11

4 Answers

vikas askiitian expert
13 years ago

net force on the system=120-50=70N

 (total mass)(accleration)=70

                  7a=70

                   a=10m/s2 ans

Chiranjeevi .P
13 years ago

a=F/m

a=(120-50)g/(1+4+2)

a=70g/7

a=10g

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

In such type of questions first of all find the net force on the system.

Here in the question,

Net force on the system=120-50=70N

 (total mass)(accleration)=70

                  7a=70

                   a=10m/s2

Hence acceleration of the system = 10 m/s2

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi



Souradeep Majumder
13 years ago

BOTH ANSWER IS CORRECT.